OpenStudy (anonymous):

my final two questions 2x+3y+3z=5 4x+5y+3z=13 use matrices and the other is find the sum of the infinite geometric series: 9,-6+4... choices for the first A.(5,13,13) b.(2,4,3) c.(1,2,-1) d.(1,3,-2)the choices for the second question are a.2/3 b.4//9 c.5 d.27/5

Directrix (directrix):

2x+3y+3z=5 4x+5y+3z=13 Is there supposed to be a third equation with these two?

OpenStudy (anonymous):

nope it only gave me two weird

Directrix (directrix):

Well, what are the instructions? I don't know what you are supposed to do with the two equations.

OpenStudy (anonymous):

oh yea forget that question use the factor theorem to determine if (x-1) is a factor of x^3+x^2-16x-16? a.factor b.not factor

Directrix (directrix):

>>oh yea forget that question Which question should be forgotten? I am totally mixed up as to whether you are asking 3 questions or what. I really wish you would post exactly one question per thread.

OpenStudy (anonymous):

never in my life have i heard of a factor theorem my teacher is retarded

OpenStudy (anonymous):

forget about this question the solving the matrices one cause i dont have the three equations

Directrix (directrix):

Take this: x^3+x^2-16x-16 and evaluate it for 1. 1^3 + 1^2 - 16 * 1 - 16 = ?

OpenStudy (anonymous):

so would the anwser be 16

Directrix (directrix):

Definition of Factor Theorem •According to Factor Theorem, if s is the root of any polynomial P(x), i.e. if P(s) = 0, then (x - s) is the factor of the polynomial P(x).

Directrix (directrix):

1^3 + 1^2 - 16 * 1 - 16 = 1 + 2 - 16 - 16 = -29

Directrix (directrix):

So, (x - 1) is not a factor of x^3+x^2-16x-16

Directrix (directrix):

find the sum of the infinite geometric series: 9,-6+4...

Directrix (directrix):

Is this supposed to be: 9 ,-6, 4... instead of 9,-6+4...

OpenStudy (anonymous):

nope thats how it was given

Directrix (directrix):

There is something wrong with this: 9,-6+4... See what it is. A series should look like this: 9 - 6 + 4 --> No commas.

OpenStudy (anonymous):

oh i must of wrote it wrong that looks write

Directrix (directrix):

I wonder if this is correct: 9 - 6 + 4 What do you think the cm is?

OpenStudy (anonymous):

these are my choices a.2/3 b.4/9 c.5 d. 27/5 i think its 5

OpenStudy (anonymous):

not sure

Directrix (directrix):

Is the series arithmetic, geometric, or something else? I wish you had the textbook so that you could check to see if you copied the problems correctly.

Directrix (directrix):

If a sequence of values follows a pattern of adding a fixed amount from one term to the next, it is referred to as an arithmetic sequence. The number added to each term is constant (always the same). The fixed amount is called the common difference, d, referring to the fact that the difference between two successive terms yields the constant value that was added. To find the common difference, subtract the first term from the second term.

OpenStudy (anonymous):

just forget it can you help me with this one and we will call it even

OpenStudy (anonymous):

im so sorry for the confusion

Directrix (directrix):

I don't see a common ratio. I don't see a common difference so the formulas for finding the sum cannot be applied.

Directrix (directrix):

Which problem are we doing now?

OpenStudy (anonymous):

forget the sequence just can you help me with this one it asks what is the system of equation that describes the graph

OpenStudy (anonymous):

and well call it done

OpenStudy (anonymous):

cause i have to sleep

Directrix (directrix):

Post it here and now. Let's do it.

Directrix (directrix):

@Karina1

OpenStudy (anonymous):

what is the equation of the following graph? A.(y+3)^2/2-(x-1)^2=1 B.(y-3)^2/4-(x-1)^2/2=1 C.(y+3)^2/4-(x-1)^2/2=1 D.(y+1)^2/4-(x-3)^2/2=1

OpenStudy (anonymous):

 i suck a t drawing